# Weak Quasistability and Rajchman measures

### v1

@@ 2. Preliminaries: Notation and Terminology
Let $\mathcal{X}$} be a normed space and let $\mathcal{X}$$^{*}\!$ be its dual. An $\mathcal{X}$}-valued sequence $\{x_{n}\}$ is weakly convergent if there exists an $x\in{\mathcal{X}}$} such that $\lim_{n}f(x_{n})=f(x)$} for every $f\in{\mathcal{X}}^{*}$}; equivalently, such that $\lim_{n}f(x_{n}\!-x)=0$} for every $f\in{\mathcal{X}}^{*}\kern-1.0pt$}. In this case the sequence $\{x_{n}\}$ is said to converge weakly to the vector $x.$ An alternative and usual notation for weak convergence that will be used here is $x_{n}\!{{\buildrel{}_{\scriptstyle w}\over{{\;\longrightarrow\;}}}}x$}\, or $x\kern-1.0pt=\kern-1.0pt{w\hbox{\,-}\lim_{n}x_{n}}.$
The above definition is standard. Actually, “every normed space has a topology $\mathcal{T}$} [the weak topology on it] such that a sequence in the space converges weakly to an element of the space if and only if the sequence converges to that element with respect to $\mathcal{T}$}. For the moment, the statement that a sequence converges weakly to a certain limit should not be taken to imply anything more than is stated [above]” [15, p.116]. We will not deal with weak topology techniques here (although the above definition coincides with convergence in the weak topology in a Hilbert space).
@@ 3. Weak Quasistability: Definition and Comparison
By an operator on a normed space $\mathcal{X}$} we mean a bounded linear (i.e., a continuous linear) transformation of $\mathcal{X}$} into itself. Let ${\mathcal{B}}[{\mathcal{X}}]$} stand for the normed algebra of all operators on $\mathcal{X}$}. The same notation $\|\;\;\|$} will be used for the norm on $\mathcal{X}$} and for the induced uniform norm on ${\mathcal{B}}[{\mathcal{X}}]$}. An operator $T$ is strongly stable if $\lim_{n}\!\|T^{n}x\|=0$} for every $x\in{\mathcal{X}}$} (notation: $T^{n}\!{{\buildrel{}_{\scriptstyle s}\over{{\;\longrightarrow\;}}}}O$}), and uniformly stable if $\lim_{n}\!\|T^{n}\|=0$} (notation: ${T^{n}\!{{\buildrel{}_{\scriptstyle u}\over{{\;\longrightarrow\;}}}}O}).$ An operator $T$ is power bounded if $\sup_{n}\!\|T^{n}\|\kern-1.0pt<\!\infty$}; equivalently, if $\sup_{n}\!\|T^{n}x\|\kern-1.0pt<\!\infty$} for every $x$ in $\mathcal{X}$} by the Banach–Steinhaus Theorem if $\mathcal{X}$} is a Banach space. If $\mathcal{X}$} is a Hilbert space, then it will be denoted by $\mathcal{H}$}, and $T^{*}\!\in{{\mathcal{B}}[{\mathcal{H}}]}$} will stand for the adjoint of $T\kern-1.0pt\in{{\mathcal{B}}[{\mathcal{H}}]}$}.
@@
Notation: $w\hbox{\,-}\lim_{n}T^{n}x=0$}, or $T^{n}x{{\buildrel{}_{\scriptstyle w}\over{{\;\longrightarrow\;}}}}0$}, for every $x\in{\mathcal{X}}$};\, or simply $T^{n}\!{{\buildrel{}_{\scriptstyle w}\over{{\;\longrightarrow\;}}}}O$}. (b) Equivalently, an operator $T$ on a normed space $\mathcal{X}$} is weakly stable if, for each vector $x\in{\mathcal{X}}$} and every subsequence $\{n_{j}\}$ of the positive integers,
@@
Notation: $w\hbox{\,-}\liminf_{n}T^{n}x=0$} for every $x\in{\mathcal{X}}$}. (b) Equivalently, an operator $T$ on a normed space $\mathcal{X}$} is weakly quasistable if for each vector $x\in{\mathcal{X}}$} there exists a subsequence $\{n_{j}\}=\{n_{j}(x)\}$ of the positive integers (that may depend on each $x$ but does not depend on $f$) for which
@@
(c) So a weakly quasistable is an operator $T$ such that, for every $x$, there is at least one subsequence $\{n_{j}\}$} (that may depend on $x$) of weak stability of $T$ for $x.$
It is worth noticing that the notion of quasistability has a plain meaning only in the weak case. In fact, it is easy to see (cf. [13, Propositions 4.1 and 4.2]) that
@@
(i.e., for power-bounded operators, quasistability coincides with stability for the strong case). For the weak case the notions are different even for power-bounded operators [13, Proposition 4.3]. However, if a weakly quasistable operator $T$ is such that for every $x$ there is a subsequence $\{n_{j}\}$ of weak stability that is of bounded increments, then $T$ is weakly stable [13, Theorem 5.3].
@@
(b) Equivalently, there is a subsequence of weak stability common to all $x\in{\mathcal{X}}$}. (c) So a homogeneously weakly quasistable is a weakly quasistable operator $T$ such that at least one subsequence $\{n_{j}\}$} of weak stability of $T$ does not depend on $x.$
@@
Notation: $T^{n}\!{{{{\buildrel{}_{\scriptstyle w}\over{{\;\longrightarrow\;}}}}\kern-13.0pt\hbox{{\fiverm/}}}\kern 9.0pt}O$}. (b) Equivalently, $T\kern-1.0pt\in{{\mathcal{B}}[{\mathcal{X}}]}$} is weakly unstable if there exists a vector $x_{0}\in{\mathcal{X}}$}, a functional $f_{0}\in{\mathcal{X}}^{*}\!$}, and a subsequence $\{m_{j}\}$ of the positive integers such that
@@
Remark 3.5. The particular case of operators acting on a Hilbert space I. If $\{T_{n}\}$ is a ${\mathcal{B}}[{\mathcal{X}}]$}-valued sequence of operators on a complex inner product space $\mathcal{X}$} with inner product $\langle\;\,\,$$\rangle$}, then the polarisation identity ensures that
@@
Such an equivalence still holds in a real Hilbert space if each $T_{n}$ is self-adjoint.
Remark 3.6. The particular case of operators acting on a Hilbert space II. The Riesz Representation Theorem in a Hilbert space ensures that if $\mathcal{H}$} is a Hilbert space with inner product $\langle\;\,\,$$\rangle$}, then according to Definitions 3.1 and 3.2, an operator $T$ on $\mathcal{H}$} is weakly stable if and only if
@@
Remark 3.7. The particular case of operators acting on a Hilbert space III. (a) Weak stability for an operator $T$ on a Hilbert space $\mathcal{H}$} means that the power sequence $\{T^{n}\}$ satisfies the equivalent limiting conditions (a) or (b) in Remark 3.6. Strong stability for an operator $T$ on a Hilbert space $\mathcal{H}$} means $\lim_{n}\|T^{n}x\|\!=0$} for every $x\in{\mathcal{H}}$}, where $\|\;\;\|$} is the norm generated by the inner product $\langle\;\,\,$$\rangle$}, which clearly implies (but is not implied by) weak stability. Differently from the strong stability case, an operator $T$ and its adjoint $T^{*}$ are weakly stable together. (b) For weak quasistability we cannot expect in general a simplified counterpart with $y\kern-1.0pt=\kern-0.5ptx$}, as in the weak stability case of Remark 3.6(b). The reason is this:: for an arbitrary weakly quasistable operator $T\kern-1.0pt$, the subsequence $\{n_{j}\}$ appearing in Remark 3.6(d) may depend on $x.$ Indeed, the polarisation identity displayed in Remark 3.5 only ensures that for each $x\in{\mathcal{H}}$} there is a subsequence $\{n_{j}(x)\}$ such that
@@ 4. Weak Quasistability and Power Boundedness
Consider the Foguel operator $F\kern-1.0pt=\kern-1.0pt\big{(}\hbox{}\,\vbox{ \halign{ \hfil${\scriptstyle#}$ \hfil&& \hfil${\scriptstyle#}$ \hfil\cr\mathstrut\crcr\kern-8.0pt\cr S^{*}&P\cr O&S^{\phantom{*}}\cr\mathstrut\crcr\kern-8.0pt\cr}}\!\big{)}$} on the direct sum $\mathcal{H}$$\oplus{\mathcal{H}}$} of a separable infinite-dimensional Hilbert space $\mathcal{H}$} with itself. Here $S$ is a unilateral shift of multiplicity one acting on $\mathcal{H}$} that shifts an orthonormal basis $\{{\bf e}_{k}\}_{k\geq 0}$ for $\mathcal{H}$}. The operator $P\!:{\mathcal{H}}\!\to{\mathcal{H}}$} is the orthogonal projection onto the closure of the span of $\{{\bf e}_{j}\!:j\in{\mathbb{J}\kern 0.5pt}\}$, where $\mathbb{J}\kern 0.5pt$} is any sparse infinite subset of positive integers with the following property: if $i,j\in{\mathbb{J}\kern 0.5pt}$} and i<j{i<j}, then 2i<j{2i<j} (e.g., ${\mathbb{J}\kern 0.5pt}={\{j\geq 1\!:\;j=3^{k};\;k\geq 0\}}$, the set of all integral powers of 3, is a sample of a sparse set of positive integers satisfying the above propriety). The operator $F$ was the first example of a power-bounded operator that is not similar to a contraction [6, 7]. Moreover, it is also known that $F$ is not weakly stable. Since, for every $n\geq 0$},
@@
It was shown in [13, Proposition 4.3] that the power-bounded and weakly unstable operator $F$ is weakly quasistable by exhibiting a subsequence $\{n_{j}\}$ of the positive integers such that each $P_{n_{j}}x_{2}$ is orthogonal to $y_{1}$ for every $x_{2},y_{1}\kern-1.0pt\in{\mathcal{H}}$}, and so $\{\langle P_{n_{j}}x_{2}\,;y_{1}\rangle\}$ is a null subsequence.
@@
(b2)  for any vector $x_{0}\kern-1.0pt\in\kern-1.0pt{\mathcal{X}}$} such that $\sup_{n}\|T^{n}x_{0}\|=\infty$}, every pair of subsequences, one of coercivity for $x_{0}$ and the other of weak stability for $x_{0}$, has a finite set of common entries.
@@
(§) $x_{n}{{\buildrel{}_{\scriptstyle w}\over{{\;\longrightarrow\;}}}}x\;\;\hbox{for some}\;\,x\in{\mathcal{X}}\;\;\Longrightarrow\;\;{\sup}_{n}\|x_{n}\|<\infty.$
Now let $\mathcal{X}$} be a Banach space and let $T$ be an operator on $\mathcal{X}$}. (a) The above implication and an application of the Banach–Steinhaus Theorem ensure the well-known relation:
@@
$T^{n_{j}}x_{0}{{\buildrel{}_{\scriptstyle w}\over{{\;\longrightarrow\;}}}}0$
@@
Remark 4.4. Power unbounded weakly quasistable operators. As is also widely known, if $T$ is a Banach-space operator, then
$r(T)\kern-1.0pt<1$ $\kern-6.0pt\iff\kern-6.0pt$ $\,T^{n}\kern-3.0pt{{\buildrel{}_{\scriptstyle u}\over{{\;\longrightarrow\;}}}}O\;\;\Longrightarrow\;\;T^{n}\kern-3.0pt{{\buildrel{}_{\scriptstyle s}\over{{\;\longrightarrow\;}}}}O\;\;\Longrightarrow\;\;T^{n}\kern-3.0pt{{\buildrel{}_{\scriptstyle w}\over{{\;\longrightarrow\;}}}}O$
@@
[…]
Also, if $T$ is a power-bounded operator on a normed space $\mathcal{X}$}, then strong quasistability coincides with strong stability [13, Proposition 4.2], that is,
@@
differently from the weak case, where there are power bounded weakly quasistable operators that are not weakly stable, as we saw in Example 4.1. However, even for the strong case, the equivalence collapses for a power unbounded operator:
${\liminf}_{n}\|T^{n}x\|\kern-1.0pt=\kern-1.0pt0\,\;\forall\,x\kern-1.0pt\in\kern-1.5pt{\mathcal{X}}\;{\kern 3.0pt\not\kern-3.0pt\Longrightarrow}\;\,{\sup}_{n}\|T^{n}\|\kern-1.0pt<\kern-1.0pt\infty\quad\&\quad{\sup}_{n}\|T^{n}\|\kern-1.0pt=\kern-1.0pt\infty\,\;{\Longrightarrow}\;\kern 0.5ptT^{n}\kern-1.0pt\!{{{{\buildrel{}_{\scriptstyle s}\over{{\;\longrightarrow\;}}}}\kern-13.0pt\hbox{{\fiverm/}}}\kern 9.0pt}O.$

### v2

# Advances in Mathematical Sciences and Applications 35(2) (2026) 559–578

@@ 2. Preliminaries: Notation and Terminology
Let $\mathcal{X}$} be a normed space and let $\mathcal{X}$$^{*}\!$ be its dual. An $\mathcal{X}$}-valued sequence $\{x_{n}\}$ is weakly convergent if there exists an $x\in{\mathcal{X}}$} such that $\lim_{n}f(x_{n})=f(x)$} for every $f\in{\mathcal{X}}^{*}$}; equivalently, such that $\lim_{n}f(x_{n}\!-x)=0$} for every $f\in{\mathcal{X}}^{*}\kern-1.0pt$}. In this case the sequence $\{x_{n}\}$ is said to converge weakly to the vector $x.$ An alternative and usual notation for weak convergence that will be used here is $x_{n}\!{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle w}}}}}x$}\, or $x\kern-1.0pt=\kern-1.0pt{w\hbox{\,-}\lim_{n}x_{n}}.$
The above definition is standard. Actually, “every normed space has a topology $\mathcal{T}$} [the weak topology on it] such that a sequence in the space converges weakly to an element of the space if and only if the sequence converges to that element with respect to $\mathcal{T}$}. For the moment, the statement that a sequence converges weakly to a certain limit should not be taken to imply anything more than is stated [above]” [16, p.116]. We will not deal with weak topology techniques here (although the above definition coincides with convergence in the weak topology in a Hilbert space).
@@ 3. Weak Quasistability: Definition and Comparison
By an operator on a normed space $\mathcal{X}$} we mean a bounded linear (i.e., a continuous linear) transformation of $\mathcal{X}$} into itself. Let ${\mathcal{B}}[{\mathcal{X}}]$} stand for the normed algebra of all operators on $\mathcal{X}$}. The same notation $\|\;\;\|$} will be used for the norm on $\mathcal{X}$} and for the induced uniform norm on ${\mathcal{B}}[{\mathcal{X}}]$}. An operator $T$ is strongly stable if $\lim_{n}\!\|T^{n}x\|=0$} for every $x\in{\mathcal{X}}$} (notation: $T^{n}\!{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle s}}}}}O$}), and uniformly stable if $\lim_{n}\!\|T^{n}\|=0$} (notation: ${T^{n}\!{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle u}}}}}O}).$ An operator $T$ is power bounded if $\sup_{n}\!\|T^{n}\|\kern-1.0pt<\!\infty$}; equivalently, if $\sup_{n}\!\|T^{n}x\|\kern-1.0pt<\!\infty$} for every $x$ in $\mathcal{X}$} by the Banach–Steinhaus Theorem if $\mathcal{X}$} is a Banach space. If $\mathcal{X}$} is a Hilbert space, then it will be denoted by $\mathcal{H}$}, and $T^{*}\!\in{{\mathcal{B}}[{\mathcal{H}}]}$} will stand for the adjoint of $T\kern-1.0pt\in{{\mathcal{B}}[{\mathcal{H}}]}$}.
@@
Notation: $w\hbox{\,-}\lim_{n}T^{n}x=0$}, or $T^{n}x{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle w}}}}}0$}, for every $x\in{\mathcal{X}}$};\, or simply $T^{n}\!{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle w}}}}}O$}.
(b) Equivalently, an operator $T$ on a normed space $\mathcal{X}$} is weakly stable if, for each vector $x\in{\mathcal{X}}$} and every subsequence $\{n_{j}\}$ of the positive integers,
@@
Notation: $w\hbox{\,-}\liminf_{n}T^{n}x=0$} for every $x\in{\mathcal{X}}$}.
(b) Equivalently, an operator $T$ on a normed space $\mathcal{X}$} is weakly quasistable if for each vector $x\in{\mathcal{X}}$} there exists a subsequence $\{n_{j}\}=\{n_{j}(x)\}$ of the positive integers (that may depend on each $x$ but does not depend on $f$) for which
@@
(c) So a weakly quasistable operator is an operator $T$ such that, for every $x$, there is at least one subsequence $\{n_{j}\}$} (that may depend on $x$) of weak stability of $T$ for $x.$
Weak quasistability plays an important role in weak l-sequential supercyclicity [13, Corollaries 4.3 and 4.4].
It is worth noticing that the notion of quasistability has a plain meaning only in the weak case. In fact, it is easy to see (cf. [14, Propositions 4.1 and 4.2]) that
@@
(i.e., for power-bounded operators, quasistability coincides with stability for the strong case). For the weak case the notions are different even for power-bounded operators [14, Proposition 4.3]. However, if a weakly quasistable operator $T$ is such that for every $x$ there is a subsequence $\{n_{j}\}$ of weak stability that is of bounded increments, then $T$ is weakly stable [14, Theorem 5.3].
@@
(b) Equivalently, there is a subsequence of weak stability common to all $x\in{\mathcal{X}}$}.
(c) So a homogeneously weakly quasistable is a weakly quasistable operator $T$ such that at least one subsequence $\{n_{j}\}$} of weak stability of $T$ does not depend on $x.$
@@
Notation: $T^{n}\!{{{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle w}}}}}\kern-13.0pt\hbox{{\fiverm/}}}\kern 9.0pt}O$}.
(b) Equivalently, $T\kern-1.0pt\in{{\mathcal{B}}[{\mathcal{X}}]}$} is weakly unstable if there exists a vector $x_{0}\in{\mathcal{X}}$}, a functional $f_{0}\in{\mathcal{X}}^{*}\!$}, and a subsequence $\{m_{j}\}$ of the positive integers such that
@@
Remark 3.5. The particular case of operators acting on a Hilbert space I.
If $\{T_{n}\}$ is a ${\mathcal{B}}[{\mathcal{X}}]$}-valued sequence of operators on a complex inner product space $\mathcal{X}$} with inner product $\langle\;\,\,$$\rangle$}, then the polarisation identity ensures that
@@
uch an equivalence still holds in a real Hilbert space if each $T_{n}$ is self-adjoint.
Remark 3.6. The particular case of operators acting on a Hilbert space II.
The Riesz Representation Theorem in a Hilbert space ensures that if $\mathcal{H}$} is a Hilbert space with inner product $\langle\;\,\,$$\rangle$}, then according to Definitions 3.1 and 3.2, an operator $T$ on $\mathcal{H}$} is weakly stable if and only if
@@
Remark 3.7. The particular case of operators acting on a Hilbert space III.
(a) Weak stability for an operator $T$ on a Hilbert space $\mathcal{H}$} means that the power sequence $\{T^{n}\}$ satisfies the equivalent limiting conditions (a) or (b) in Remark 3.6. Strong stability for an operator $T$ on a Hilbert space $\mathcal{H}$} means $\lim_{n}\|T^{n}x\|\!=0$} for every $x\in{\mathcal{H}}$}, where $\|\;\;\|$} is the norm generated by the inner product $\langle\;\,\,$$\rangle$}, which clearly implies (but is not implied by) weak stability. Differently from the strong stability case, an operator $T$ and its adjoint $T^{*}$ are weakly stable together.
(b) For weak quasistability we cannot expect in general a simplified counterpart with $y\kern-1.0pt=\kern-0.5ptx$}, as in the weak stability case of Remark 3.6(b). The reason is this:: for an arbitrary weakly quasistable operator $T\kern-1.0pt$, the subsequence $\{n_{j}\}$ appearing in Remark 3.6(d) may depend on $x.$ Indeed, the polarisation identity displayed in Remark 3.5 only ensures that for each $x\in{\mathcal{H}}$} there is a subsequence $\{n_{j}(x)\}$ such that
@@ 4. Weak Quasistability and Power Boundedness
Consider the Foguel operator $F\kern-1.0pt=\kern-1.0pt\big(\hbox{}\,\vbox{\halign{ \hfil${\scriptstyle#}$ \hfil&& \hfil${\scriptstyle#}$ \hfil\cr\mathstrut\crcr\kern-8.0pt\cr S^{*}&P\cr O&S^{\phantom{*}}\cr\mathstrut\crcr\kern-8.0pt\cr}}\!\big)$} on the direct sum $\mathcal{H}$$\oplus{\mathcal{H}}$} of a separable infinite-dimensional Hilbert space $\mathcal{H}$} with itself. Here $S$ is a unilateral shift of multiplicity one acting on $\mathcal{H}$} that shifts an orthonormal basis $\{{\bf e}_{k}\}_{k\geq 0}$ for $\mathcal{H}$}. The operator $P\!:{\mathcal{H}}\!\to{\mathcal{H}}$} is the orthogonal projection onto the closure of the span of $\{{\bf e}_{j}\!:j\in{\mathbb{J}\kern 0.5pt}\}$, where $\mathbb{J}\kern 0.5pt$} is any sparse infinite subset of positive integers with the following property: if $i,j\in{\mathbb{J}\kern 0.5pt}$} and i<j{i<j}, then 2i<j{2i<j} (e.g., ${\mathbb{J}\kern 0.5pt}={\{j\geq 1\!:\;j=3^{k};\;k\geq 0\}}$, the set of all integral powers of 3, is a sample of a sparse set of positive integers satisfying the above property). The operator $F$ was the first example of a power-bounded operator that is not similar to a contraction [6, 7]. Moreover, it is also known that $F$ is not weakly stable. Since, for every $n\geq 0$},
@@
It was shown in [14, Proposition 4.3] that the power-bounded and weakly unstable operator $F$ is weakly quasistable by exhibiting a subsequence $\{n_{j}\}$ of the positive integers such that each $P_{n_{j}}x_{2}$ is orthogonal to $y_{1}$ for every $x_{2},y_{1}\kern-1.0pt\in{\mathcal{H}}$}, and so $\{\langle P_{n_{j}}x_{2}\,;y_{1}\rangle\}$ is a null subsequence.
@@
(b2)   for any vector $x_{0}\in{\mathcal{X}}$}\, such that $\sup_{n}\|T^{n}x_{0}\|=\infty$},\, every pair of subsequences, one of coercivity for $x_{0}$ and the other of weak stability for $x_{0}$, has a finite set of common entries.
@@
(§) $x_{n}{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle w}}}}}x\;\;\hbox{for some}\;\,x\in{\mathcal{X}}\;\;\Longrightarrow\;\;{\sup}_{n}\|x_{n}\|<\infty.$
Now let $\mathcal{X}$} be a Banach space and let $T$ be an operator on $\mathcal{X}$}.
(a) The above implication and an application of the Banach–Steinhaus Theorem ensure the well-known relation:
@@
$T^{n_{j}}x_{0}{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle w}}}}}0$
@@
Remark 4.4. Power unbounded weakly quasistable operators.
As is also widely known, if $T$ is a Banach-space operator, then
$r(T)\kern-1.0pt<1$ $\kern-6.0pt\iff\kern-6.0pt$ $\,T^{n}\kern-3.0pt{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle u}}}}}O\;\;\Longrightarrow\;\;T^{n}\kern-3.0pt{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle s}}}}}O\;\;\Longrightarrow\;\;T^{n}\kern-3.0pt{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle w}}}}}O$
@@
$\kern-3.0pt\Longrightarrow\kern-3.0pt$ ${\sup}_{n}\|T^{n}x\|<\infty\;\;\forall\,x\in{\mathcal{X}}.{\phantom{\Big|}}$
Also, if $T$ is a power-bounded operator on a normed space $\mathcal{X}$}, then strong quasistability coincides with strong stability [14, Proposition 4.2], that is,
@@
differently from the weak case, where there are power-bounded weakly quasistable operators that are not weakly stable, as we saw in Example 4.1. However, even for the strong case, the equivalence collapses for a power unbounded operator:
${\liminf}_{n}\|T^{n}x\|\kern-1.0pt=\kern-1.0pt0\,\;\forall\,x\kern-1.0pt\in\kern-1.5pt{\mathcal{X}}\;{\kern 3.0pt\not\kern-3.0pt\Longrightarrow}\;\,{\sup}_{n}\|T^{n}\|\kern-1.0pt<\kern-1.0pt\infty\quad\&\quad{\sup}_{n}\|T^{n}\|\kern-1.0pt=\kern-1.0pt\infty\,\;{\Longrightarrow}\;\kern 0.5ptT^{n}\kern-1.0pt\!{{{{\mathrel{\mathop{\kern 0.0pt{\;\longrightarrow\;}}\limits^{{}_{\scriptstyle s}}}}}\kern-13.0pt\hbox{{\fiverm/}}}\kern 9.0pt}O.$